OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 3 points) [7.02] Choose the correct simplification of (8x + 7) + (2x − 6). 10x − 1 10x + 13 10x + 1 11x -------------------------------------------------------------------------------- Question 3 (Multiple Choice Worth 3 points) [7.03] Choose the correct simplification of (x9)2. x11 x7 x81 x18 -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 4 points) [7.03] Choose the correct simplification of 4x0. 4 4x 1 0 -------------------------------------

OpenStudy (anonymous):

1) 8x + 7 + 2x − 6 = 10x +1 2) x^18 3) 0

OpenStudy (anonymous):

I'm a bit doubtful about the last question -- is it 4*0 or 4x^0 ?

OpenStudy (anonymous):

its 4^0

OpenStudy (anonymous):

my bad 4x^0

OpenStudy (anonymous):

K .. so 4x^0 = 4 * 1 = 4 .. Since , a^0 = 1 (Law of exponents)

OpenStudy (anonymous):

thanks i never knew there was a law of exponets.

OpenStudy (anonymous):

but wouldn't it be like 4x? or did the x get canceled out?

OpenStudy (anonymous):

no, it's just a law.. anything to the power 0 , becomes 1 .. i.e.. a^0 = 1 there's one more law that anything to the power one is the same thing . i.e a^1 = a .. Hope you got your answer :)

OpenStudy (anonymous):

i did and thamk you

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

You're welcome